# PocketGS: On-Device Training of 3D Gaussian Splatting for High Perceptual Modeling

### v2

@@ I Introduction
Training 3DGS on device introduces substantial challenges, as it must deliver high perceptual quality under the limited computational resources and strict thermal constraints of mobile platforms. Consequently, the training process must meet on-device constraints: 1) completion within a few minutes ( $\textless 5$ min), 2) respecting the hardware-available memory budget ( $\textless 3$ GB).
These constraints require rethinking 3DGS under resource-limited regimes beyond conventional resource-unconstrained assumptions. Gaussian-based modeling typically includes: (1) recovering camera parameters and a point cloud from images, (2) initializing 3D Gaussians from this geometry, and (3) optimizing the Gaussians via differentiable rendering for scene recovery.
Existing designs assume ample time and memory across all stages, but on-device settings break these assumptions, yielding three tightly coupled contradictions.
@@
Hardware-Differentiability Contradiction. On-device 3DGS is at odds with mobile execution: differentiable rendering requires storing intermediates for backpropagation, but mobile pipelines restrict gradient access through opaque blending, suffer from costly GPU–CPU synchronization in optimization, and incur heavy data movement and state misalignment from depth-ordered alpha blending. Therefore, on-device 3DGS requires a mobile-oriented differentiable rendering framework that bounds and replays only the necessary intermediates for Gaussian optimization.
In this work, we present PocketGS, a 3DGS training paradigm that achieves high-perceptual modeling under on-device constraints via three co-designed operators (Fig. LABEL:top,2). To address the Input-Recovery Contradiction, the geometry-prior construction operator $\mathcal{G}$ builds a geometry-faithful prior, providing the required geometric guidance without costly in-loop Gaussian densification beyond the on-device envelope. To resolve the Initialization-Convergence Contradiction, the prior-conditioned Gaussian parameterization operator $\mathcal{I}$ uses this prior to shape initial Gaussians, improving conditioning for stable, iteration-limited optimization and reducing degradation from noise-sensitive initialization. Finally, to overcome the Hardware-Differentiability Contradiction, the hardware-aligned splatting optimization operator $\mathcal{T}$ makes differentiable rendering practical on mobile GPUs, preserving correct gradients and stable updates without prohibitive memory traffic or CPU–GPU synchronization. Extensive experiments show that PocketGS achieves perceptual quality superior to mainstream workstation 3DGS baselines, while operating within on-device constraints.
@@
  • We validate that, while satisfying on-device constraints, PocketGS achieves competitive perceptual quality relative to standard workstation 3DGS pipelines on publicly available benchmarks and our self-collected MobileScan datasets.
Figure 2: Overview of the PocketGS framework. PocketGS tackles on-device 3DGS training through three coupled operators: ($\mathcal{G}$) Geometry Prior Construction employs an information-gated gate for frame selection, followed by GPU-native Schur-complement BA and single-reference MVS. The MVS module constructs a 3D cost volume by sampling depth hypotheses $d_{1},\dots,d_{n}$} via census matching to produce a dense geometric scaffold. ($\mathcal{I}$) Prior-Conditioned Parameterization seeds anisotropic Gaussians by estimating local surface statistics (normals $\mathbf{s}_{n}$ and scales) to front-load structure discovery via disc-like covariance seeding. ($\mathcal{T}$) Hardware-Aligned Splatting implements a mobile-native differentiable renderer using unrolled alpha-compositing ($S={C_{in},\alpha}$) and index-mapped gradient scattering to ensure stable backpropagation within tight mobile memory bounds of the canonical parameter buffer ($\mu,\Sigma,c,\alpha$).
@@ III The PocketGS Paradigm
PocketGS enables on-device 3DGS training: given a short mobile capture $I_{t}$}, it optimizes a Gaussian scene $\Theta$ under strict runtime and peak-memory budgets on tile-based mobile GPUs. It co-designs three coupled operators: $\mathcal{G}$ derives a compact geometry prior $\mathcal{P}$ from $I_{t}$}, $\mathcal{I}$ uses $\mathcal{P}$ to initialize well-conditioned Gaussians $\Theta_{0}$, and $\mathcal{T}$ performs hardware-aligned differentiable splatting to keep training GPU-resident with correct gradients. Formally, $\mathcal{G}:(\{I_{t}\},\{\hat{T}_{t}\})\mapsto(\{T_{t}\},\mathcal{P}),\ \mathcal{I}:\mathcal{P}\mapsto\Theta_{0},$ and $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$, balancing perceptual quality with on-device constraints.
@@ III-A Geometry Prior Construction under Resource Budgets
To resolve the Input-Recovery Contradiction, Operator $\mathcal{G}$ constructs a low-noise and low-redundant 3DGS-aligned prior within mobile compute and memory constraints. Treating BA (Bundle Adjustment) and MVS (Multi-View Stereo) as core prior-construction primitives, their outputs directly condition the 3DGS optimization, mitigating over-reliance on excessive in-loop Gaussian densification.
@@ III-A1 Information-Gated Frame Subsampling
To reduce input noise and provide high-quality photometric ground truth for 3DGS training, PocketGS adopts a real-time three-stage gating mechanism that admits only frames with incremental geometric utility, maximizing information yield per compute unit.
Displacement Gate. We ensure sufficient parallax for robust triangulation by computing the camera displacement $d$ between the current frame and the last selected keyframe, $d=\|\mathbf{t}_{curr}-\mathbf{t}_{last}\|_{2}$. A frame is considered only if $d\geq\tau_{d}$, where $\tau_{d}=0.05m$, suppressing redundant viewpoints while maintaining tracking stability.
@@
$S=\frac{1}{|\Omega|}\sum_{(x,y)\in\Omega}\left(|I(x+\Delta,y)-I(x,y)|+|I(x,y+\Delta)-I(x,y)|\right)$ (1)
@@
Candidate Windowing. We select the locally optimal frame from a short temporal window (8 frames or 250ms). A candidate replaces the current best only if $S_{new}>(1+r)S_{best}$, with $r=0.05$. This commits the sharpest representative per viewpoint and bounds downstream optimization workload.
Keyframe admission is explicitly bounded by the three gates, reducing downstream BA/MVS workload roughly proportional to admitted keyframes rather than raw capture length, thereby constraining both runtime and peak intermediate storage.
@@ III-A2 GPU-Native Global BA as Mobile MAP Refinement
To address mobile input imprecision, PocketGS implements GPU-native global BA to jointly optimize poses and 3D points, delivering 3DGS-required precision without over-densification. The robustified reprojection objective for PocketGS:
$\min_{\{\mathbf{T}_{i}\},\{\mathbf{P}_{j}\}}\sum_{i,j}\rho\left(\|\pi(\mathbf{T}_{i},\mathbf{P}_{j})-\mathbf{p}_{ij}\|^{2}_{\Sigma_{ij}}\right)$ (2)
where $\pi$ is the projection function, $\mathbf{T}_{i}\in SE(3)$ is the pose of camera $i$, $\mathbf{P}_{j}\in\mathbb{R}^{3}$ is the 3D position of point $j$, $\mathbf{p}_{ij}\in\mathbb{R}^{2}$ is the observed 2D projection, and $\rho$ is the robust Huber loss.
@@
GPU-Native Schur Complement. The BA bottleneck is solving the normal equations $\mathbf{H}\Delta=\mathbf{b}$, where $\mathbf{H}=\mathbf{J}^{T}\mathbf{J}$. We exploit the sparse block structure via a full GPU Schur complement. Partitioning parameters into camera pose (c) and point (p) blocks yields the reduced camera system:
$(\mathbf{H}_{cc}-\mathbf{H}_{cp}\mathbf{H}_{pp}^{-1}\mathbf{H}_{pc})\Delta_{c}=\mathbf{b}_{c}-\mathbf{H}_{cp}\mathbf{H}_{pp}^{-1}\mathbf{b}_{p}$ (3)
Since $\mathbf{H}_{pp}$ is block-diagonal with independent $3\times 3$ blocks per point, $\mathbf{H}_{pp}^{-1}$ is computed in parallel across points on the GPU. After solving for $\Delta_{c}$, we recover $\Delta_{p}$ by back-substitution.
@@
Schur complement reduces solving from the full joint system to a smaller camera system, while $\mathbf{H}_{pp}^{-1}$ is computed in parallel; the design keeps memory dominated by sparse observations rather than dense matrices, making global refinement feasible on-device.
@@ III-A3 Single-Reference Cost-Volume MVS
To address mobile input sparsity, PocketGS’s lightweight MVS generates a dense geometry-faithful point cloud, providing structural richness for Gaussian initialization without over-densification.
@@ III-B Prior-Conditioned Gaussian Parameterization
To ensure stable convergence under a strict mobile iteration budget, operator $\mathcal{I}$ improves early-stage conditioning by replacing isotropic initialization with prior-conditioned, surface-aligned anisotropic Gaussians, thereby injecting structural cues from $\mathcal{G}$ and reducing reliance on photometric signals to infer geometry.
@@ III-B1 Local Surface Statistics Estimation
To inject geometric prior knowledge into initialization, we estimate local surface statistics for each point $\mathbf{p}_{i}$ in the dense MVS point cloud via KNN (K=16). The neighborhood covariance matrix is computed as:
$\mathbf{C}=\frac{1}{K}\sum_{k=1}^{K}(\mathbf{p}_{k}-\bar{\mathbf{p}})(\mathbf{p}_{k}-\bar{\mathbf{p}})^{T}$ (5)
where $\bar{\mathbf{p}}_{i}$ is the centroid of the neighborhood.The eigenvector corresponding to $\mathbf{C}_{i}$’s smallest eigenvalue robustly estimates the local surface normal $\mathbf{n}_{i}$. This GPU-parallelized process provides critical structural cues for informed initialization.
Fixed local neighborhood size ensures embarrassingly parallel computation with bounded per-point memory, aligning with PocketGS’s mobile constraints.
III-B2 Disc-Like Covariance Seeding
We initialize each Gaussian as a thin, disc-like ellipsoid tangent to the estimated surface, directly addressing the contradiction by replacing naive isotropic initialization with geometry-conditioned anisotropy. The tangential scale $s_{t}$ is derived from local point density, estimated as the average distance to the K=3 nearest neighbors:
$s_{t}=\frac{1}{3}\sum_{k=1}^{3}\|\mathbf{p}-\mathbf{p}_{k}\|$ (6)
We set the normal-direction scale $s_{n}$ as a fraction of tangential scale:
$s_{n}=s_{t}\times r_{normal}$ (7)
where $r_{normal}=0.3$. Scale parameters are optimized in log-space, $\mathbf{s}_{log}=\log(\mathbf{s})$, ensuring positivity and numerical stability. The initial rotation $\mathbf{q}$ aligns the Gaussian local z-axis with the estimated surface normal $\mathbf{n}_{i}$. Opacity is initialized to a fixed low value (logit of 0.1). This geometry-conditioned anisotropy supplies strong priors on surface orientation and extent, accelerating convergence under limited iterations.
@@ III-C Hardware-Aligned Differentiable Splatting Optimization
Operator $\mathcal{T}$ rewrites the 3DGS training loop for GPU-resident optimization on mobile GPUs under tight memory/bandwidth budgets, reducing intermediate storage and synchronization overhead while enabling differentiable blending. By unrolling compositing and decoupling parameter/state layout from rendering order, it supports end-to-end on-device training with correct, stable gradients.
@@ III-C1 Unrolled Alpha-Compositing with Forward Replay Cache
To address hidden intermediate states, PocketGS unrolls alpha-compositing into an explicit differentiable pipeline with cached intermediates. Standard front-to-back compositing follows:
@@
where $C_{in}$ is the pre-composite color, and $(\alpha,c)$ are the opacity and color of the current Gaussian fragment. In a mobile rendering pipeline, hardware blending returns only $C_{out}$ while hiding intermediate factors, making correct gradient propagation unavailable without costly framebuffer reads.
Explicit Computation Graph.
We process visible Gaussians in depth-sorted batches. Each batch explicitly takes the previous accumulator as input, exposing the intermediates necessary for the backward pass:
$\frac{\partial\mathcal{L}}{\partial C_{in}}=\frac{\partial\mathcal{L}}{\partial C_{out}}(1-\alpha),\quad\frac{\partial\mathcal{L}}{\partial\alpha}=\frac{\partial\mathcal{L}}{\partial C_{out}}(c-C_{in}).$ (9)
Bandwidth-Efficient Cache.
To minimize memory bandwidth pressure-a critical constraint for mobile energy efficiency-we introduce a forward replay cache. Instead of caching full per-pixel splat lists, we store a minimal set $S=\{C_{in},\alpha\}$ and use a per-pixel counter buffer. By resetting only the $O(WH)$ counter rather than the entire $O(WHK_{max})$ cache, we significantly reduce the bandwidth overhead during the backward traversal.

### v3

@@ I Introduction
Training 3DGS on device introduces substantial challenges, as it must deliver high perceptual quality under the limited computational resources and strict thermal constraints of mobile platforms. Consequently, the training process must meet on-device constraints: 1) completion within a few minutes (e.g., <5<5 min), 2) respecting the hardware-available memory budget (e.g., <3<3 GB).
These constraints require rethinking 3DGS under resource-limited regimes beyond conventional resource-unconstrained assumptions. Gaussian-based modeling typically includes: (1) recovering camera parameters and a point cloud from images, (2) initializing 3D Gaussians from this geometry, and (3) optimizing the Gaussians via differentiable rendering for scene recovery. Existing designs assume ample time and memory across all stages, but on-device settings break these assumptions, yielding three tightly coupled contradictions. 21
@@
Hardware–Differentiability Mismatch. On-device 3DGS conflicts with mobile execution: while differentiable splatting needs per-step alpha-compositing states for backpropagation, mobile fixed-function pipelines expose only the final blended color and conceal the depth-ordered accumulation needed to route gradients. Recovering these states via framebuffer readbacks, full buffer clears, or backward-time re-enumeration of per-pixel splat sequences is bandwidth-prohibitive and poorly aligned with mobile GPU state. Thus, on-device 3DGS calls for a mobile-oriented differentiable rendering framework that explicitly unrolls compositing and replays only the minimal cached intermediates required for Gaussian optimization.
In this work, we present PocketGS, a 3DGS training paradigm that achieves high-perceptual modeling under on-device constraints via three co-designed operators (Fig. 1,2). To address the Input-Recovery Contradiction, the geometry-prior construction operator $\mathcal{G}$ builds a geometry-faithful prior, providing the required geometric guidance without costly in-loop Gaussian densification beyond the on-device envelope. To resolve the Initialization-Convergence Contradiction, the prior-conditioned Gaussian parameterization operator $\mathcal{I}$ uses this prior to shape initial Gaussians, improving conditioning for stable, iteration-limited optimization and reducing degradation from noise-sensitive initialization. Finally, to overcome the Hardware-Differentiability Contradiction, the hardware-aligned splatting optimization operator $\mathcal{T}$ makes differentiable rendering practical on mobile GPUs, preserving correct gradients and stable updates without prohibitive memory traffic. Extensive experiments show that PocketGS achieves perceptual quality superior to mainstream workstation 3DGS baselines, while operating within on-device constraints.
@@
We empirically show that PocketGS satisfies strict on-device runtime and peak-memory budgets while achieving competitive perceptual quality compared with standard workstation 3DGS pipelines, on both publicly available benchmarks and our self-collected MobileScan dataset.
Figure 2: Overview of the PocketGS framework. PocketGS tackles on-device 3DGS training through three coupled operators: ($\mathcal{G}$) Geometry Prior Construction employs an information-gated gate for frame selection, followed by GPU-native Schur-complement BA and single-reference MVS. The MVS module constructs a 3D cost volume by sampling depth hypotheses $d_{1},\dots,d_{n}$} via census matching to produce a dense geometric scaffold. ($\mathcal{I}$) Prior-Conditioned Parameterization seeds anisotropic Gaussians by estimating local surface statistics (normals $\mathbf{s}_{n}$ and scales) to front-load structure discovery via disc-like covariance seeding. ($\mathcal{T}$) Hardware-Aligned Splatting implements a mobile-native differentiable renderer using unrolled alpha-compositing ($S={id,C_{in},\alpha}$) and index-mapped gradient scattering to ensure stable backpropagation within tight mobile memory bounds of the canonical parameter buffer ($\mu,\Sigma,c,\alpha$).
@@ III The PocketGS Paradigm
PocketGS enables on-device 3D Gaussian Splatting (3DGS) training under strict runtime and peak-memory constraints. Given a sequence of video frames $\{I_{t}\}_{t=1}^{n}$ and their corresponding coarse camera poses $\{\hat{T}_{t}\}_{t=1}^{n}$ obtained from mobile tracking systems (ARKit or ARCore), PocketGS comprises three tightly coupled operators. Specifically, the geometry operator $\mathcal{G}$ refines the coarse poses $\{\hat{T}_{t}\}$ into accurate ones $\{{T}_{t}\}$, and extracts a dense point cloud $P$, which serves as a reliable geometric prior. The initialization operator $\mathcal{I}$ maps $P$ to a well-conditioned 3D Gaussian initialization $\Theta_{0}$. Finally, the training operator $\mathcal{T}$ performs hardware-aligned differentiable rendering to optimize $\Theta_{0}$ into the final Gaussian representation $\Theta^{*}$. Formally, these operators are defined as: $\mathcal{G}:(\{I_{t},\hat{T}_{t}\})\mapsto(\{T_{t}\},P)$, $\mathcal{I}:P\mapsto\Theta_{0}$, and $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$.
@@ III-A Geometry Prior Construction under Resource Budgets
Prior 3DGS approaches rely on prolonged optimization with training-time densification to recover scenes from noisy and sparse point inputs. On mobile devices, this paradigm often yields unstable geometry recovery and substantial computational overhead, making it impractical under on-device constraints. We instead leverage the mobile capture $\{I_{t},\hat{T}_{t}\}$ to build a dense, low-noise, and low-redundancy geometry prior by applying a geometry operator $\mathcal{G}$, which maps $(\{I_{t},\hat{T}_{t}\})$ to refined poses $\{T_{t}\}$ and a compact geometric representation $P$. Conditioning the subsequent optimization on $(\{T_{t}\},P)$ removes the need for training-time densification, resulting in more stable convergence and improved computational efficiency.
@@ III-A1 Information-Gated Frame Subsampling
Video frames captured during mobile scanning are often affected by motion blur, which degrades geometric reconstruction and undermines the photometric optimization of 3DGS. To mitigate this issue, we subsample the input sequence using an information-gated keyframe selection strategy based on viewpoint change and distance variation, retaining only sharp and high-utility frames while filtering out blurred and redundant ones.
Displacement Gate. We ensure sufficient parallax for robust triangulation by computing the camera displacement $d$ between the current frame $\mathbf{t}_{curr}$ and the last selected keyframe $\mathbf{t}_{last}$, $d=\|\mathbf{t}_{curr}-\mathbf{t}_{last}\|_{2}$. A frame is considered only if $d\geq\tau_{d}$, where $\tau_{d}=0.05m$, suppressing redundant viewpoints while maintaining tracking stability.
@@
$\small S=\frac{1}{|\Omega|}\sum_{(x,y)\in\Omega}\left(|I(x+\Delta,y)-I(x,y)|+|I(x,y+\Delta)-I(x,y)|\right)$ (1)
@@
Candidate Windowing. We select the locally optimal frame from a short temporal window (8 frames). A candidate replaces the current best only if $S_{new}>(1+r)S_{best}$, with $r=0.05$. This commits the sharpest representative per viewpoint and bounds downstream optimization workload.
@@ III-A2 GPU-Native Global BA as Mobile MAP Refinement
Initial camera poses provided by mobile tracking systems (e.g., ARKit and ARCore) are often noisy, which degrades both geometric reconstruction and the photometric optimization of 3DGS. To address this issue, we perform GPU-native global BA on-device to refine poses, thereby providing accurate pose estimates for subsequent dense point recovery and 3DGS optimization.
The robustified reprojection objective for PocketGS:
$\min_{\{\mathbf{\hat{T}}_{i}\},\{\mathbf{P}_{j}\}}\sum_{i,j}\rho\left(\|\pi(\mathbf{\hat{T}}_{i},\mathbf{P}_{j})-\mathbf{p}_{ij}\|^{2}_{\Sigma_{ij}}\right)$ (2)
where $\pi$ is the projection function, $\mathbf{\hat{T}}_{i}\in SE(3)$ is the pose of camera $i$, $\mathbf{P}_{j}\in\mathbb{R}^{3}$ is the 3D position of point $j$, $\mathbf{p}_{ij}\in\mathbb{R}^{2}$ is the observed 2D projection, and $\rho$ is the robust Huber loss.
@@
GPU-Native Schur Complement. The BA bottleneck is solving the normal equations $\mathbf{H}\Delta=\mathbf{b}$, where $\mathbf{H}=\mathbf{J}^{T}\mathbf{J}$. We exploit the sparse block structure via a full GPU Schur complement. Partitioning parameters into camera pose (t) and point (p) blocks yields the reduced camera system:
$(\mathbf{H}_{tt}-\mathbf{H}_{tp}\mathbf{H}_{pp}^{-1}\mathbf{H}_{pt})\Delta_{c}=\mathbf{b}_{t}-\mathbf{H}_{tp}\mathbf{H}_{pp}^{-1}\mathbf{b}_{p}$ (3)
Since $\mathbf{H}_{pp}$ is block-diagonal with independent $3\times 3$ blocks per point, $\mathbf{H}_{pp}^{-1}$ is computed in parallel across points on the GPU. After solving for $\Delta_{t}$, we recover $\Delta_{p}$ by back-substitution.
@@ III-A3 Single-Reference Cost-Volume MVS
Point clouds provided by mobile tracking systems are typically sparse, which necessitates costly densification during 3DGS training. To address this limitation, our lightweight MVS reconstructs a dense and geometry-faithful point cloud for Gaussian initialization, thereby reducing over-densification under mobile budgets.
@@ III-B Prior-Conditioned Gaussian Parameterization
Prior methods initialize points as isotropic Gaussians based solely on inter-point distances, resulting in arbitrary parameter initialization that requires unnecessary updates and consequently slow convergence. In contrast, we exploit point-cloud surface geometry $P$ to estimate normals and initialize surface-aligned anisotropic Gaussians via an initialization operator $\mathcal{I}$, which maps $P$ to the initial Gaussian parameters $\Theta_{0}$. This structure-consistent starting state $\Theta_{0}=\mathcal{I}(P)$ improves conditioning and accelerates optimization under tight on-device budgets.
@@ III-B1 Local Surface Statistics Estimation
To extract geometric priors from the mobile capture, we estimate local surface statistics for each point $\mathbf{p}_{i}$ in the dense MVS point cloud using a fixed-size KNN neighborhood (K=16K{=}16). Let $\{\mathbf{p}_{i,k}\}_{k=1}^{K}$ denote the $K$ nearest neighbors of $\mathbf{p}_{i}$, with centroid
$\bar{\mathbf{p}}_{i}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{p}_{i,k}.$ (5)
We compute the local covariance as
$\mathbf{C}_{i}=\frac{1}{K}\sum_{k=1}^{K}(\mathbf{p}_{i,k}-\bar{\mathbf{p}}_{i})(\mathbf{p}_{i,k}-\bar{\mathbf{p}}_{i})^{\mathsf{T}}.$ (6)
The eigenvector of $\mathbf{C}_{i}$ associated with its smallest eigenvalue provides an estimate of the local surface normal $\mathbf{n}_{i}$. This per-point, fixed-$K$ procedure is embarrassingly parallel with bounded memory, making it well-suited for GPU execution under PocketGS’s mobile constraints.
@@ III-C Hardware-Aligned Differentiable Splatting Optimization
On mobile GPUs, on-device 3D Gaussian Splatting (3DGS) training is bandwidth-bound: fixed-function blending reveals only the final color, hiding the intermediate compositing states needed for backpropagation, while auxiliary-buffer clears incur heavy memory traffic. We therefore align optimization with the mobile pipeline by unrolling front-to-back compositing into an explicit differentiable operator and caching a compact forward replay trace. This defines a GPU-resident map $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$ that delivers correct, stable gradients without framebuffer readbacks or backward-time reconstruction of per-pixel splat sequences.
@@ III-C1 Unrolled Alpha-Compositing with Forward Replay Cache
Mobile fixed-function blending returns only the final composited color, while 3DGS backpropagation requires the intermediate accumulator state at each compositing step. Moreover, the set and order of Gaussians contributing to a pixel are not accessible from the rendering pipeline, making a backward pass that replays the original per-pixel splat list infeasible or prohibitively expensive. We address this by (1) unrolling the front-to-back compositing into an explicit computation graph and (2) caching a compact forward replay trace per pixel for backward traversal.
Explicit Computation Graph.
For depth-sorted visible Gaussians, we explicitly perform front-to-back compositing:
@@
where $C_{in}$ is the accumulator input and $(\alpha,c)$ are the opacity and color of the current Gaussian fragment. The backward pass follows the same unrolled chain:
$\frac{\partial\mathcal{L}}{\partial C_{in}}=\frac{\partial\mathcal{L}}{\partial C_{out}}(1-\alpha),\qquad\frac{\partial\mathcal{L}}{\partial\alpha}=\left\langle\frac{\partial\mathcal{L}}{\partial C_{out}},\,(c-C_{in})\right\rangle,$ (8)
where $\langle\cdot,\cdot\rangle$ denotes the per-channel inner product.
Bandwidth-Efficient Forward Replay Cache.
To avoid backward-time reconstruction of per-pixel splat sequences, we record only the intermediates necessary to replay gradients:
$S=\{\mathrm{id},\,C_{in},\,\alpha\},$ (9)
where $\mathrm{id}$ is the (sorted) splat identifier used to scatter gradients back to canonical parameter buffers. A per-pixel counter $\mathrm{count}(u)$ stores the number of valid cached entries for pixel $u$. Each iteration resets only the $O(WH)$ counters, rather than clearing the full $O(WHK_{\max})$ cache. During backpropagation, we traverse only
$k\in[0,\mathrm{count}(u)-1],$
treating stale cache entries beyond $\mathrm{count}(u)$ as invalid, thereby eliminating bandwidth-heavy full-cache clears while preserving gradient correctness.